You hear on the local news that for the city of Kalamazoo, the proportion of people who support President Trump is 0.59. However, you think it is different from 0.59. The hypotheses for this test are Null Hypothesis: p = 0.59, Alternative Hypothesis: p ≠ 0.59. If you randomly sample 28 people and 10 of them support President Trump, what is your test statistic and p-value?
Answer #1

Given that, n = 28 and x = 10

point estimate = x/n = 10/28 = 0.3571

The null and the alternative hypotheses are,

Test statistic is,

Test statistic = -2.51

since, it is two-tailed test,

p-value = 2 * P(Z < -2.51) = 2 * 0.0060 = 0.0120

=> p-value = 0.0120